Lifespans of male mice have a mean of 860 days and standard deviation of 40 days. The distribution is slightly left skewed. Sample mean lifespan for a sample of 64 mice has what mean and standard deviation of days?
Answer #1

By the central limit theorem, the distribution of sample mean will be approximately normal with:

Mean days = Population mean = 860 days

Standard deviation = 40/V64 = 40/8 = 5 days